A soccer ball of mass 0.50 kg, initially at rest, is kicked by a force of magnitude 1200 N that lasts for 15 milliseconds. 

What is the velocity of the ball after the kick? 


1
Expert's answer
2021-03-17T16:56:12-0400
"F\Delta t=m\Delta v,""F\Delta t=m(v_f-v_i)=mv_f,""v_f=\dfrac{F\Delta t}{m}=\dfrac{1200\ N\cdot15\cdot10^{-3}\ s}{0.5\ kg}=36\ \dfrac{m}{s}."